# Making Differential Privacy Easier to Use for Data Controllers using a Privacy Risk Indicator

### v4

@@ 5. Releasing Epsilon Privately and Bounding Privacy Loss
In Section 5.1, we propose a modification of the Find-$\epsilon$-from-RDR algorithm that permits finding and releasing $\epsilon$ while preserving DP. In Section 5.2, we discuss how the new algorithms answer multiple queries without the need to set a total privacy budget and we provide the end-to-end DP guarantee in Section 5.3.
@@ 5.1. Finding and Releasing $\epsilon$ using SVT
Key Idea. The Find-$\epsilon$-from-RDR algorithm tests each candidate $\epsilon$ in descending order and chooses the largest $\epsilon$ that passes the threshold ($\tau_{p}$) specified by the controller. Our key idea is to make the threshold differentially private: we leverage the Sparse Vector Technique (SVT) (Dwork et al., 2009) to test candidate $\epsilon$ and release the first $\epsilon$ that approximately passes the threshold, so releasing it preserves DP.
@@
1
[…]
3for $q_{i}\in\mathcal{Q}_{svt}$ do
4
5      if $q_{i}(x)+Lap\left(\frac{2\Delta_{svt}}{\epsilon_{svt}^{\prime\prime}}\right)\geq\tau+\rho$ then
6             Output $a_{i}=\top$
7             return
8      else
9             Output $a_{i}=\bot$
10
@@ 9             Output $a_{i}=\bot$
What is a good SVT query for choosing $\epsilon$? Intuitively, each SVT query in our modified algorithm will be computing $RDR_{min}/RDR_{max}$ for each candidate $\epsilon$. Then the algorithm applies SVT by comparing $RDR_{min}/RDR_{max}$ to $\tau_{p}$ and releasing $\epsilon$ such that $RDR_{min}/RDR_{max}$ is approximately greater than or equal to $\tau_{p}$. However, this approach has one caveat: the SVT query has a sensitivity $\Delta_{svt}=1$ since a single change of record in the dataset can cause the corresponding $RDR_{min}/RDR_{max}$ to change by at most 1. This means to apply SVT, the algorithm will add a Laplacian noise scaled to $1/\epsilon_{svt}^{\prime\prime}$ to $\tau_{p}$ and another Laplacian noise scaled to $2/\epsilon_{svt}^{\prime\prime}$ to $RDR_{min}/RDR_{max}$. Clearly, when $\epsilon_{svt}$ is small (e.g., less than 1), the threshold will be too noisy and make the algorithm useless for finding $\epsilon$.
To address this caveat, we opt for a different SVT query that computes the variance of normalized RDRs. For each candidate $\epsilon$, the algorithm computes the variance of the within-dataset individuals’ RDRs normalized between 0 and 1, and compares the variance to a threshold ($\tau_{var}$) specified by the controller. If the variance is approximately smaller than $\tau_{var}$, then a suitable $\epsilon$ is found. The SVT query has a sensitivity of $\Delta_{svt}=1/n$, where n is the number of records in the dataset; this means when $n$ is relatively large, the Laplacian noise that is added to the variance and $\tau_{var}$) will be small.
Choosing threshold $\tau_{var}$. Although it is intuitive to interpret the variance of RDRs — small variance means the within-dataset individuals’ disclosure risk are similar and large variance means some within-dataset individuals are at higher risk than others, setting an appropriate threshold $\tau_{var}$ is non-trivial. Consider the example we presented in Section 4.1 where the query is to count the number of patients who have a certain disease, but instead of a dataset of three records, we have a large dataset with only one patient who has the disease. Then the variance of the patients’ RDRs will be small for any $\epsilon$ since all but one patient have the same RDR. In this case, the controller needs to choose $\tau_{var}$ that is extremely close to or exactly 0 to protect the patient who has the disease. In general, we recommend that controllers first observe the computed per-instance sensitivity of each within-dataset individual before choosing an appropriate $\tau_{var}$.
@@
1
[…]
3
[…]
5
6/* Generate noise to add to threshold */
7 $\rho=Lap\left(\frac{\Delta_{svt}}{\epsilon_{svt}^{\prime}}\right)$
8
[…]
10
[…]
12
13for $x_{i}\in x^{\prime}$ do
14       if $x_{i}\notin PIS$ then
15             $PIS[x_{i}]\leftarrow\textsc{compute-per-instance-sensitivity}(x^{\prime},x_{i},q,M)$
16
17
18/* Controllers may observe the computed per-instance sensitivity of each individual first and set $\tau_{var}$ based on their privacy preferences */
19
20for $\epsilon\in\mathcal{E}$ do
21
22      $RDR\leftarrow list()$
23       for $x_{i}\in x^{\prime}$ do
24             $RDR_{i}\leftarrow\textsc{compute-rdr}(PIS[x_{i}],M,q,\epsilon,\delta)$
25             $RDR.add(RDR_{i})$
26
27      $\textsc{normalize}(RDR)$
28       $Var\leftarrow$ compute-variance($RDR$)
29
30      /* Run SVT to find $\epsilon$ that approximately passes threshold */
31
32      if $-Var+Lap\left(\frac{2\Delta_{svt}}{\epsilon_{svt}^{\prime\prime}}\right)\geq-\tau_{var}+\rho$ then
33
34            $o\leftarrow\mathcal{M}(x^{\prime},q,\epsilon,\delta)$
35             return $\epsilon,o$
36
@@ 34            $o\leftarrow\mathcal{M}(x^{\prime},q,\epsilon,\delta)$
The Find-and-release-$\epsilon$-from-RDR algorithm. We present the Find-and-release-$\epsilon$-from-RDR algorithm (Algorithm 3). The main modification lies in line 3 -  3 and line 3 -  3. In line 3, the algorithm allocate the SVT budget $\epsilon_{svt}$; one allocation strategy (Lyu et al., 2016) is to have $\epsilon_{svt}^{\prime}:\epsilon_{svt}^{\prime\prime}=1:2^{2/3}$ to achieve better accuracy of SVT. Then it iterates over each candidate $\epsilon$ and computes its corresponding RDRs the same way as in Algorithm 1 (line 3 -  3). It then normalizes each RDR (e.g., dividing by the maximum RDR) and computes the variance (line 3 - line 3). Lastly, it compares the noisy variance with the noisy threshold (controllers want $Var\leq\tau_{var}$, which is equivalent to $-Var\geq-\tau_{var}$). If the variance passes the threshold (line 3), the algorithm returns $\epsilon$ and DP output $o$ (line 3).
The main advantage of using SVT is that it always consumes a fixed budget $\epsilon_{svt}$ regardless of the number of SVT queries (i.e., the number of candidate $\epsilon$ values). Controllers understand the consequence of running SVT since they know exactly how much noise is added to both the variance and the threshold $\tau_{var}$. They can decide whether they are comfortable choosing an $\epsilon$ that does not exactly conform to their privacy preferences; if not, they may use a different $\epsilon_{svt}$ or $\tau_{var}$ and rerun the algorithm.
@@ 5.2. Bounding Privacy Loss Across Queries
Considering analysts’ preferences over queries. A natural tension exists between analysts and controllers: analysts want more accurate output, whereas controllers want to protect the within-dataset individuals’ privacy by releasing noisy output. Because it gets increasingly harder to find a suitable $\epsilon$ after answering multiple queries as $\epsilon_{c}$ keeps increasing, analysts may need to prioritize queries of which they need more accurate output. One approach to incorporate analysts’ preferences when finding $\epsilon$ is to allow them to assign weights (i.e., priority scores) to their queries, then controllers can adjust the parameter $\tau_{p}$ or $\tau_{var}$ in each algorithm based on analysts’ preferences, e.g., they can set a smaller $\tau_{p}$ or larger $\tau_{var}$ for high-priority queries so the algorithm can find larger $\epsilon$. Nevertheless, solving related problems, such as how to communicate analysts’ preferences to controllers, and how to help controllers decide an appropriate $\tau_{p}$ or $\tau_{var}$ that takes both analysts’ preferences and their own privacy preferences into account, are outside the scope of the paper but remain interesting directions to explore.
@@ 6.1.1. Recruitment and data collection
We recruited 100 participants using Prolific (pro, [n.d.]). To complete the study, each participant must sign a consent form, which describes the purpose and content of the study, compensation, and data confidentiality. We do not collect identifiable information, and the survey responses are anonymized, meaning the participant’s IP address, location data, and contact information are not being recorded.
@@ 6.1.2. Study procedure
First, the participants are randomly divided into control or treatment groups; 29 participants are in the control group, and 27 are in the treatment group. The survey presents a sample from the UCI Adult dataset (uci, [n.d.]) and explains a DP mechanism is used to answer queries on the dataset by adding noise controlled by the parameter $\epsilon$. Then, it presents two queries in random order:
@@ 6.1.4. Threats to validity
6.2. RQ2: Are Both Algorithms Practical?
Datasets. We controlled the experiment by generating datasets with 1000, 10000, 100000, and 1 million records from the Adult dataset (uci, [n.d.]). We sampled records for datasets of sizes (number of records) 1000 and 10000. For datasets of size 100000 and 1 million, we replicated and added random noise to the replicated records to ensure that they were not the same copy of each other. We included all 15 attributes in the dataset.
@@ 6.2. RQ2: Are Both Algorithms Practical?
Parameters. We initialize candidate $\epsilon$ values from 11 to 0.001 in fixed intervals; specifically, we test $\epsilon$ in $[10,\cdots,1,0.9,\cdots,0.1,0.09,\\ \cdots,0.01,0.099,\cdots,0.001]$ (37 values in total). We set $\tau_{p}=0.95$ in Find-$\epsilon$-from-RDR, and $\epsilon_{svt}=1,\tau_{var}=10^{-5}$ in Find-and-release
-$\epsilon$-from-RDR
. We used Laplace Mechanism as the underlying DP mechanism $\mathcal{M}$. We used 8 parallel processes to compute the per-instance sensitivity of each within-dataset individual.
Configurations. We ran the experiments on a Chameleon Cloud (cha, [n.d.]) instance (Intel Xeon Gold 6242 CPU) with 32 cores and 188GB memory. We used Ubuntu 24.04 and Python 3.12 for all experiments.
@@ 6.3.1. Effect of $\tau_{p}$ in Find-$\epsilon$-from-RDR
The algorithm finds $\epsilon$ by letting controllers specify their privacy preferences through a threshold $\tau_{p}$, and checks whether $RDR_{min}/RDR_{max}\geq\tau_{p}$ for each candidate $\epsilon$. Larger $\tau_{p}$ means stricter privacy preferences (thus smaller $\epsilon$) because $RDR_{min}$ and $RDR_{max}$ need to be closer. We test the effect of varying $\tau_{p}$ in the algorithm and report the $\epsilon$ found.
@@ 6.3.2. Effect of $\tau_{var}$ in Find-and-release-$\epsilon$-from-RDR
The algorithm lets controllers specify the variance threshold $\tau_{var}$ (and the SVT budget $\epsilon_{svt}$), and applies SVT to choose the first $\epsilon$ such that the variance of normalized RDRs is approximately smaller than $\tau_{var}$. Smaller $\tau_{var}$ means stricter privacy preferences. Because of the randomness introduced by SVT, the algorithm may find different $\epsilon$ with the same input parameters. We evaluate the effect of varying $\tau_{var}$ by running the algorithm 50 times for each parameter setting and report $\epsilon$ found in each round.

### v5

@@ 5. Releasing Epsilon Privately and Bounding Privacy Loss Across Queries
In Section 5.1, we propose a modification of the Find-$\epsilon$-from-RDR algorithm that permits finding and releasing $\epsilon$ while preserving DP. In Section 5.2, we discuss how the new algorithms answer multiple queries without the need to set a total privacy budget, and we provide the end-to-end DP guarantee in Section 5.3.
@@ 5.1. Finding and Releasing $\epsilon$ using SVT
Key Idea. The Find-$\epsilon$-from-RDR algorithm iterates through candidate $\epsilon$ values and selects the first one satisfying the controller’s privacy preference, i.e., $f(RDR)\geq\tau$. Our key idea is to make this selection process itself differentially private. We leverage the Sparse Vector Technique (SVT) (Dwork et al., 2009) to test candidate $\epsilon$ values and release the first value such that $f(RDR)$ approximately passes the threshold, thereby ensuring releasing $\epsilon$ preserves DP.
@@
1ex
$\rho=Lap\left(\frac{\Delta_{svt}}{\epsilon_{svt}^{\prime}}\right)$
for $q_{i}\in\mathcal{Q}_{svt}$ do
 if $q_{i}(x)+Lap\left(\frac{2\Delta_{svt}}{\epsilon_{svt}^{\prime\prime}}\right)\geq\tau+\rho$ then
     Output $a_{i}=\top$
     return
  end if
 else
     Output $a_{i}=\bot$
  end if
end for
@@
A Low-Sensitivity SVT Query. Intuitively, the controller’s preference function becomes the SVT query. While the Find-$\epsilon$-from-RDR algorithm offers the controller the flexibility to define any preference functions, many functions have high global sensitivity, making them unsuitable for SVT. For example, the function $RDR_{min}/RDR_{max}$ has a sensitivity of 1. This means to apply SVT, the algorithm will add a Laplacian noise scaled to $1/\epsilon_{svt}^{\prime}$ to $\tau_{p}$ and another Laplacian noise scaled to $2/\epsilon_{svt}^{\prime\prime}$ to $RDR_{min}/RDR_{max}$. Clearly, when $\epsilon_{svt}$ is small (e.g., less than 1), the threshold will be too noisy and make the algorithm useless for finding $\epsilon$.
To address this caveat, we propose an SVT query based on the variance of normalized RDRs. We first normalize each within-dataset individual’s RDR, $RDR_{i}$, by a global, data-independent RDR upper bound, $RDR_{max}^{*}$, which depends only on the query’s global sensitivity, output dimensionality $k$, and the candidate $\epsilon$ value. Based on Lemmas 3.6 and 3.7, $RDR_{max}^{*}=\Delta_{1}+k\Delta_{1}/\epsilon$ if Laplace Mechanism is used, and $RDR_{max}^{*}=\sqrt{\Delta_{2}^{2}+k\sigma^{2}}$ if Gaussian Mechanism is used. Then the normalized RDR, $RDR_{i}^{\prime}$, for a candidate $\epsilon$ is $RDR_{i}^{\prime}(\epsilon)=RDR_{i}(\epsilon)/RDR_{max}^{*}(\epsilon)$.
Definition 0 (SVT query).
The SVT query for each candidate $\epsilon$ is defined as:
$q_{svt}(\epsilon)=Var(\{RDR_{i}^{\prime}(\epsilon)|i\in[n]\}),$
where $n$ is the number of within-dataset individuals.
The SVT query has a global sensitivity $\Delta_{svt}=1/n$, since $0\leq RDR_{i}^{\prime}(\epsilon)\leq 1$. This means the noise added by SVT will be very small for reasonably large datasets.
@@
1ex
$\epsilon_{svt}^{\prime},\epsilon_{svt}^{\prime\prime}\leftarrow\textsc{assign-svt-budget}(\epsilon_{svt})$
$\rho=Lap\left(\frac{\Delta_{svt}}{\epsilon_{svt}^{\prime}}\right)$
$x^{\prime}\leftarrow\textsc{project-query-attributes}(x,q)$
$PIS\leftarrow map()$
for $x_{i}\in x^{\prime}$ do
  if $x_{i}\notin PIS$ then
     $PIS[x_{i}]\leftarrow\textsc{compute-per-instance-sensitivity}(x^{\prime},x_{i},q,M)$
  end if
end for
/* Controllers may observe the computed per-instance sensitivity of each individual first, and set $\tau_{var}$ based on their privacy preferences */
for $\epsilon\in\mathcal{E}$ do
 $RDR\leftarrow list()$
  for $x_{i}\in x^{\prime}$ do
     $RDR_{i}\leftarrow\textsc{compute-rdr}(PIS[x_{i}],M,q,\epsilon,\delta)$
     $RDR.add(RDR_{i})$
  end for
 /* Run SVT to find the first $\epsilon$ that approximately passes threshold */
 if $q_{svt}(\epsilon)+Lap\left(\frac{2\Delta_{svt}}{\epsilon_{svt}^{\prime\prime}}\right)\geq-\tau_{var}+\rho$ then
    $o\leftarrow\mathcal{M}(x^{\prime},q,\epsilon,\delta)$
     return $\epsilon,o$
  end if
end for
@@
The Find-and-release-$\epsilon$-from-RDR algorithm. Algorithm 3 integrates this SVT-based query. First, the algorithm allocate the SVT budget $\epsilon_{svt}$ ( line 3); one allocation strategy (Lyu et al., 2016) is to have $\epsilon_{svt}^{\prime}:\epsilon_{svt}^{\prime\prime}=1:2^{2/3}$ to achieve better accuracy of SVT. Then it iterates over each candidate $\epsilon$ and computes the corresponding RDRs the same way as in Algorithm 1 (line 3 -  3). For each $\epsilon$, it computes the SVT query and tests if the variance is approximately below the threshold $\tau_{var}$ (line 3). The first $\epsilon$ that passes this noisy test is selected and released along with the final DP output $o$ (line 3).
The main advantage of using SVT is that it always consumes a fixed budget $\epsilon_{svt}$ regardless of the number of SVT queries (i.e., the number of candidate $\epsilon$ values). Controllers understand the consequence of running SVT since they know exactly how much noise is added to both the SVT query and the threshold. They can decide whether they are comfortable choosing an $\epsilon$ that does not exactly conform to their privacy preferences; if not, they may use a different $\epsilon_{svt}$ or $\tau_{var}$ and rerun the algorithm.
@@
Claim: Assume the SVT budget $\epsilon_{svt}$ budget is split such that $\epsilon_{svt}^{\prime}=\epsilon_{svt}/3$ and $\epsilon_{svt}^{\prime}=2\epsilon_{svt}/3$. For any failure probability $\beta\in(0,1)$, Find-and-release-$\epsilon$-from-RDR algorithm is $(\alpha,\beta)$-accurate, where $\alpha=\frac{6}{n\epsilon_{svt}}\ln(\frac{2|\mathcal{E}|}{\beta})$; the accuracy of the algorithm refers to the correctness of the chosen $\epsilon$. This means that with probability at least $1-\beta$, if the algorithm returns a value $\epsilon^{*}$, then the true variance associated with that value satisfies $q_{svt}(\epsilon^{*})\leq\tau_{var}+\alpha$. And for any candidate $\epsilon^{\prime}\in\mathcal{E}$ such that $q_{svt}(\epsilon^{\prime})<\tau_{var}-\alpha$, the algorithm will not halt at a value $\epsilon<\epsilon^{\prime}$, thus returning a value at least as large as $\epsilon^{\prime}$(since candidate $\epsilon$ values are tested in descending order). A complete proof is provided in the Appendix.
@@ 5.2. Bounding Privacy Loss Across Queries
Considering analysts’ preferences over queries. A natural tension exists between analysts and controllers: analysts want more accurate output, whereas controllers want to protect the within-dataset individuals’ privacy by releasing noisy output. Because it gets increasingly harder to find a suitable $\epsilon$ after answering multiple queries as $\epsilon_{c}$ keeps increasing, analysts may need to prioritize queries of which they need more accurate output. One approach to incorporate analysts’ preferences when finding $\epsilon$ is to allow them to assign weights (i.e., priority scores) to their queries, then controllers can adjust the threshold $\tau$ based on analysts’ preferences, e.g., they can set a smaller $\tau_{p}$ or larger $\tau_{var}$ for high-priority queries so the algorithm can find larger $\epsilon$. Nevertheless, solving related problems, such as how to communicate analysts’ preferences to controllers, and how to help controllers decide an appropriate $\tau$ that takes both analysts’ accuracy preferences and their own privacy preferences into account, are outside the scope of the paper but remain interesting directions to explore.
@@ 6.1.1. Recruitment and data collection
We recruited 100 participants using Prolific (34). To complete the study, each participant must sign a consent form, which describes the purpose and content of the study, compensation, and data confidentiality. We do not collect identifiable information, and the survey responses are anonymized, meaning the participant’s IP address, location data, and contact information are not being recorded.
@@ 6.1.2. Study procedure
First, the participants are randomly divided into control or treatment groups; 29 participants are in the control group, and 27 are in the treatment group. The survey presents a sample from the UCI Adult dataset (40) and explains a DP mechanism is used to answer queries on the dataset by adding noise controlled by the parameter $\epsilon$. Then, it presents two queries in random order:
@@ 6.2. RQ2: Are Both Algorithms Practical?
Datasets. We controlled the experiment by generating datasets with 1000, 10000, 100000, and 1 million records from the Adult dataset (40). We sampled records for datasets of sizes (number of records) 1000 and 10000. For datasets of size 100000 and 1 million, we replicated and added random noise to the replicated records to ensure that they were not the same copy of each other. We included all 15 attributes in the dataset.
@@
Parameters. We initialize candidate $\epsilon$ values in fixed intervals; specifically, we test $\epsilon$ in $10,\cdots,1,0.9,\cdots,0.1,0.09,\cdots,0.01,0.099,\\ \cdots,0.001]$ (37 values in total). We use $RDR_{min}/RDR_{max}$ as the preference function and set $\tau_{p}=0.95$ in Find-$\epsilon$-from-RDR, and we set $\epsilon_{svt}=1,\tau_{var}=10^{-5}$ in Find-and-release-$\epsilon$-from-RDR. We used Laplace Mechanism as the underlying DP mechanism $\mathcal{M}$. We used 8 parallel processes to compute the per-instance sensitivity of each within-dataset individual.
Configurations. We ran the experiments on a Chameleon Cloud (4) instance (Intel Xeon Gold 6242 CPU) with 32 cores and 188GB memory. We used Ubuntu 24.04 and Python 3.12 for all experiments.
@@ 6.3.1. Effect of $\tau_{p}$ in Find-$\epsilon$-from-RDR
Given the preference function $RDR_{min}/RDR_{max}$, we test the effect of varying threshold $\tau_{p}$ and report the $\epsilon$ found.
@@ 6.3.2. Effect of $\tau_{var}$ in Find-and-release-$\epsilon$-from-RDR
Due to the randomness of SVT, we evaluate the effect of varying $\tau_{var}$ by running the algorithm 50 times and report $\epsilon$ found in each round.
@@ 8. Conclusions
Appendix A Appendix A: Derivation of RDR Upper Bounds (Lemmas 3.1 and 3.2)
Lemma A.1.
$RDR_{i}$ upper bound under Laplace Mechanism is
$\lVert q(x)-q(x_{-i})\rVert_{1}+k\Delta_{1}/\epsilon.$
Proof.
Let $o=q(x)+Z$, where $Z=(Z_{1},\dots,Z_{k})$ and $Z_{j}\sim Lap(b)$ are i.i.d. with $b=\Delta_{1}/\epsilon$. Let $C_{i}=q(x)-q(x_{-i})$.
By Definition 1 and 5 in the paper, $RDR_{i}=\mathbb{E}_{o\sim\mathcal{M}_{Lap}(x,q,\epsilon)}\left[\|o-q(x_{-i})\|_{1}\right]=\mathbb{E}[\|C_{i}+Z\|_{1}]$.
By the definition of the $\ell_{1}$ norm and the linearity of expectation: $\mathbb{E}[\|C_{i}+Z\|_{1}]=\mathbb{E}\left[\sum\limits_{j=1}^{k}|(C_{i})_{j}+Z_{j}|\right]=\sum\limits_{j=1}^{k}\mathbb{E}[|(C_{i})_{j}+Z_{j}|]$.
By the triangle inequality: $\sum\limits_{j=1}^{k}\mathbb{E}[|(C_{i})_{j}+Z_{j}|]\leq\sum\limits_{j=1}^{k}\mathbb{E}[|(C_{i})_{j}|+|Z_{j}|]$.
Since $(C_{i})_{j}$ is a constant for each dimension $j$, this simplifies to: $\sum\limits_{j=1}^{k}\mathbb{E}[|(C_{i})_{j}|+|Z_{j}|]=\left(\sum\limits_{j=1}^{k}|(C_{i})_{j}|\right)+\left(\sum\limits_{j=1}^{k}\mathbb{E}[|Z_{j}|]\right)=\|C_{i}\|_{1}+\left(\sum\limits_{j=1}^{k}\mathbb{E}[|Z_{j}|]\right)$.
Since $Z_{j}\sim Lap(b)$, $\mathbb{E}[|Z_{j}|]=b=\Delta_{1}/\epsilon$. Since all $Z_{j}$ are i.i.d., $\sum\limits_{j=1}^{k}\mathbb{E}[|Z_{j}|]=\sum\limits_{j=1}^{k}k\Delta_{1}/\epsilon$.
Combining the terms and substituting back $C_{i}=q(x)-q(x_{-i})$, we get the upper bound: $RDR_{i}\leq\lVert q(x)-q(x_{-i})\rVert_{1}+k\Delta_{1}/\epsilon$.
This completes the proof.
Lemma A.2.
$RDR_{i}$ upper bound under Gaussian Mechanism is
$\sqrt{\lVert q(x)-q(x_{-i})\rVert_{2}^{2}+k\sigma^{2}}.$
Proof.
Let $o=q(x)+Z$, where $Z=(Z_{1},\dots,Z_{k})$ and $Z_{j}\sim\mathcal{N}(0,\sigma^{2})$ are i.i.d. with $\sigma^{2}=\frac{2\Delta_{2}^{2}\ln(1.25/\delta)}{\epsilon^{2}}$. Let $C_{i}=q(x)-q(x_{-i})$.
By Definition 1 and 5 in the paper, $RDR_{i}=\mathbb{E}_{o\sim\mathcal{M}_{Gau}(x,q,\epsilon,\delta)}[\|o-q(x_{-i})\|_{2}]=\mathbb{E}[\|C_{i}+Z\|_{2}]$.
By Jensen’s inequality, $(RDR_{i})^{2}=(\mathbb{E}[\|C_{i}+Z\|_{2}])^{2}\leq\mathbb{E}[\|C_{i}+Z\|_{2}^{2}]$.
By linearity of expectation: $\mathbb{E}[\|C_{i}+Z\|_{2}^{2}]=\mathbb{E}\left[\sum\limits_{j=1}^{k}((C_{i})_{j}+Z_{j})^{2}\right]=\sum\limits_{j=1}^{k}\mathbb{E}[((C_{i})_{j}+Z_{j})^{2}]$.
Expanding the squared term: $\sum\limits_{j=1}^{k}\mathbb{E}[(C_{i})_{j}^{2}+2(C_{i})_{j}Z_{j}+Z_{j}^{2}]=\sum\limits_{j=1}^{k}(\mathbb{E}[(C_{i})_{j}^{2}]+2(C_{i})_{j}\mathbb{E}[Z_{j}]+\mathbb{E}[Z_{j}^{2}])$.
Since $(C_{i})_{j}$ is a constant for each dimension $j$, $\mathbb{E}[(C_{i})_{j}^{2}]=(C_{i})_{j}^{2}$.
Since each $Z_{j}\sim\mathcal{N}(0,\sigma^{2})$, $\mathbb{E}[Z_{j}]=0$ and the variance $Var(Z_{j})=\mathbb{E}[Z_{j}^{2}]-(\mathbb{E}[Z_{j}])^{2}=\sigma^{2}$, thus $\mathbb{E}[Z_{j}^{2}]=\sigma^{2}$.
Substituting these values back, we get: $\mathbb{E}[\|C_{i}+Z\|_{2}]=\sum\limits_{j=1}^{k}((C_{i})_{j}^{2}+2(C_{i})_{j}\cdot 0+\sigma^{2})=\sum\limits_{j=1}^{k}\left((C_{i})_{j}^{2}\right)+\sum\limits_{j=1}^{k}\left(\sigma^{2}\right)=\|C_{i}\|_{2}^{2}+k\sigma^{2}$.
We have thus shown that: $(RDR_{i})^{2}\leq\|C_{i}\|_{2}^{2}+k\sigma^{2}$.
Taking the square root of both sides and substituting back $C_{i}=q(x)-q(x_{-i})$, we get the upper bound: $\sqrt{\lVert q(x)-q(x_{-i})\rVert_{2}^{2}+k\sigma^{2}}$.
This completes the proof.
Appendix B Appendix B: Accuracy Analysis of Find-and-release-$\epsilon$-from-RDR algorithm
Claim: Assume the SVT budget $\epsilon_{svt}$ budget is split such that $\epsilon_{svt}^{\prime}=\epsilon_{svt}/3$ and $\epsilon_{svt}^{\prime}=2\epsilon_{svt}/3$. For any failure probability $\beta\in(0,1)$, Find-and-release-$\epsilon$-from-RDR algorithm is $(\alpha,\beta)$-accurate, where $\alpha=\frac{6}{n\epsilon_{svt}}\ln(\frac{2|\mathcal{E}|}{\beta})$; the accuracy of the algorithm refers to the correctness of the chosen $\epsilon$. This means that with probability at least $1-\beta$, if the algorithm returns a value $\epsilon^{*}$, then the true variance for that value satisfies $q_{svt}(\epsilon)\leq\tau_{var}+\alpha$. And for any candidate $\epsilon^{\prime}\in\mathcal{E}$ such that $q_{svt}(\epsilon^{\prime})<\tau_{var}-\alpha$, the algorithm will not halt at a value $\epsilon<\epsilon^{\prime}$, thus returning a value at least as large as $\epsilon^{\prime}$(since candidate $\epsilon$ values are tested in descending order).
Proof.
The SVT mechanism in Find-and-release-$\epsilon$-from-RDR algorithm checks if $q_{svt}(\epsilon)+Lap\left(\frac{2\Delta_{svt}}{\epsilon_{svt}^{\prime\prime}}\right)\geq-\tau_{var}+\rho$, where $\rho=Lap\left(\frac{\Delta_{svt}}{\epsilon_{svt}^{\prime}}\right)$. Let $\nu=Lap\left(\frac{2\Delta_{svt}}{\epsilon_{svt}^{\prime\prime}}\right)$. Since we assume a budget split of $\epsilon_{svt}^{\prime}=\epsilon_{svt}/3$ and $\epsilon_{svt}^{\prime}=2\epsilon_{svt}/3$, we have $\nu\sim Lap(3\Delta_{svt}/\epsilon_{svt})$ and $\rho\sim Lap(3\Delta_{svt}/\epsilon_{svt})$; let $b=3\Delta_{svt}/\epsilon_{svt}$.
Let the total noise on the comparison be $Z=\nu-\rho$. An error occurs if the noise flips the outcome of the comparison, which happens if $|Z|>|q_{svt}(\epsilon)-\tau_{var}|$. We bound the probability that $|Z|>\alpha$ for any of the $\mathcal{E}$ queries.
Using a standard tail bound for the difference of two i.i.d. Laplace variables, $\nu,\rho$: $\Pr[|Z|>\alpha]\leq 2\exp(-\frac{\alpha}{2b})$.
Then we apply a union bound over all $\mathcal{E}$ candidate $\epsilon$ values. The probability that the noise for any query exceeds $\alpha$ is at most: $\Pr[\exists\epsilon\in\mathcal{E},|Z_{\epsilon}|>\alpha]\leq|\mathcal{E}|\cdot 2\exp(-\frac{\alpha}{2b})$.
We set this failure probability to be at most $\beta$: $|\mathcal{E}|\cdot 2\exp(-\frac{\alpha}{2b})\leq\beta\Rightarrow\alpha\geq 2b\ln(\frac{2|\mathcal{E}|}{\beta})$.
Substituting $b=3\Delta_{svt}/\epsilon_{svt}$ and $\Delta_{svt}=1/n$: $\alpha\geq 2(\frac{3(1/n)}{\epsilon_{svt}})(\ln(\frac{2|\mathcal{E}|}{\beta}))=\frac{6}{n\epsilon_{svt}}\ln(\frac{2|\mathcal{E}|}{\beta})$.
This establishes the $(\alpha,\beta)$-accuracy guarantee. With probability at least $1-\beta$, for all $|\mathcal{E}|$ queries, the noisy comparison is within $\alpha$ of the true comparison.